OpenStudy (anonymous):

Two taps can together fill a tank in 9 3/8hrs. The tap of larger diameter takes 10 hours less than the smaller one to fill the tank separately. Find the time in which each tap can separately fill the tank. I know that the equation is 1/x-10 + 1/x =8/75 but how do u get it as 1/x. Why and how? Pls explain thanks.

OpenStudy (anonymous):

1/(x-10)+1/x=8/75 multiply both side of the equation (x-10)(x)

OpenStudy (anonymous):

then the equation x+(x-10)=8/75(x)(x-10)

OpenStudy (anonymous):

How?